OpenStudy (anonymous):

find the exact value of sin(pi/12) , cos(pi/12) , and tan(pi/12) using half angle formula.

OpenStudy (jamesj):

Well, \[ \cos(2x) = 2\cos^2 x - 1 \] and you know the value of \( \cos(\pi/6) \) (right? I hope you do). Hence you can find \[ \cos(\pi/12) \] From that you can find \[ \sin(\pi/12) \] and tan.